# SCHEME: Scalable Channer Mixer for Vision Transformers

### v1

@@ A.2 More results of removing CCA at inference
Table 13 and 14 show the results of removing CCA at inference for object detection and semantic segmentation models, respectively. For both tasks, results are identical to the model using CCA showing that the result also generalizes to downstream tasks.
A.3 CCA
In the paper, feature groups across different layers of the model was used to demonstrate the learning of feature clusters by CCA. Here, the feature from the final layer of the model is only considered. Table 11 shows the top-1 accuracy per feature group and model. The average of the outputs from the four group classifiers is reported in the final column of the table. The effect is more pronounced when using a single layer feature with more than 1% accuracy difference between the model with and without CCA. This further reinforces that by introducing inter-group communication, CCA enables the groups to learn more diverse sets of features, that complement each other.
A.4 Larger expansion ratios
Table 12 shows additional results of using larger expansion ratios with SCHEME for the same parameters and FLOPs using the Metaformer-PPAA-S12 baseline model. We observe that SCHEMEformer-PPAA consistently outperforms the baseline for larger expansions ratios and saturates as the model size and FLOPs increases.
Appendix B Implementation Details
B.1 Hyperparameter Settings
Table 15 shows the detailed hyperparameter settings of the family of SCHEMEformer models reported in the main paper.
Appendix C Qualitative Analysis: Grad-CAM [38] Visualization
Fig. 6 shows the results of class activation maps for SCHEMEformer-PPAA-44-e8-S12 model for few examples from the validation set of ImageNet-1K dataset. The stronger heatmap responses around the salient features of an object (e.g., body of a bird, cat) shows that the model ignores the background and attends to more discriminative spatial regions. Fig. 6 also shows the comparison with a few existing methods that demonstrates that SCHEMEformer-PPAA-44-e8-S12 attends to the complete object class and is better than the competing methods.
(a) Input Images
(b) ResNet-50 [21]
(c) ResMLP [42]
(d) DeiT-Small[43]
(e) Poolformer-S12[57]
(f) SCHEMEformer-PPAA-44-e8-S12
@@ Appendix C Qualitative Analysis: Grad-CAM [38] Visualization
Figure 6: GRAD-CAM [38] visualization for a few validation samples on ImageNet-1K dataset for SCHEMEformer-PPAA-44-e8-S12 model.

### v4

@@ A.3.4 Impact of removing CCA at inference
Table A.3.3 and A.3.3 show the results of removing CCA at inference for object detection and semantic segmentation models, respectively. For both tasks, the results are identical to the model using CCA showing that the CCA also generalizes to downstream tasks.
Table 16: Ablation study on the effect of larger expansion ratios in BD-MLP block of SCHEME on ImageNet-1K validation dataset.
#Par FLOPs Top-1
Model (M) (G) Acc (%)
Metaformer-PPAA-11-e1-S12 9.6 1.37 76.0
SCHEMEformer-PPAA-22-e2-S12 9.6 1.37 77.0
SCHEMEformer-PPAA-44-e4-S12 9.6 1.37 77.4
SCHEMEformer-PPAA-66-e6-S12 10.0 1.45 77.9
Metaformer-PPAA-11-e2-S12 11.8 1.77 78.9
SCHEMEformer-PPAA-22-e4-S12 11.8 1.77 79.4
SCHEMEformer-PPAA-44-e8-S12 11.8 1.77 79.7
SCHEMEformer-PPAA-33-e6-S12 12.5 1.87 79.8
SCHEMEformer-PPAA-22-e6-S12 14.1 2.16 80.4
Metaformer-PPAA-11-e4-S12 16.5 2.56 81.0
SCHEMEformer-PPAA-22-e8-S12 16.5 2.56 81.1
SCHEMEformer-PPAA-44-e16-S12 16.5 2.56 81.2
Table 17: Ablation study on the formation of feature clusters in the BD-MLP branch of the SCHEME module. We train a linear classifier on top of the four feature groups extracted from the final MLP mixer of the transformer block of the network. The model trained with CCA forms feature clusters that learn diverse and complementary set of features that can obtain 1% higher validation accuracy than the model trained without CCA.
CCA Group1 Group2 Group3 Group4 Ensemble
Model (%) (%) (%) (%) (%)
SCHEMEformer-44-e8-S12 29.63 37.59 37.57 35.89 73.95
SCHEMEformer-44-e8-S12 27.89 37.64 30.02 37.67 75.04
@@
Figure 7: GRAD-CAM [33] visualization for a few validation samples on ImageNet-1K dataset for SCHEMEformer-PPAA-44-e8-S12 model and comparison with other competing methods.
A.3.5 CCA
In the main paper, feature groups across different layers of the model was used to demonstrate the learning of feature clusters by CCA. Here, the feature from the final layer of the model is only considered. Table A.3.4 shows the top-1 accuracy per feature group and model. The average of the outputs from the four group classifiers is reported in the final column of the table. The effect is more pronounced when using a single layer feature with +1.09% accuracy difference between the model with and without CCA. This further reinforces that by introducing inter-group communication, CCA enables the groups to learn more diverse sets of features, that complement each other.
A.3.6 Larger expansion ratios
Table A.3.4 shows additional results of using larger expansion ratios with SCHEME mixer (illustrated in Fig. 3 of the main paper) for the same number of parameters and FLOPs using the Metaformer-PPAA-S12 baseline model. We observe that SCHEMEformer-PPAA consistently outperforms the baseline for larger expansions ratios with larger gains at lower FLOPs. The performance saturates as the model size and FLOPs increases.
Figure 8: Visualization of last layer attention maps of the SCHEMEformer-PPAA-44-e8-S12 model with and without CCA on Imagenet validation dataset images. The ellipses identify the region of the object in the target class. Since the SCHEMEformer model does not have class tokens, the attention maps are not necessarily interpretable. The spikes in the attention maps are also common in attention visualizations and can be avoided by using registers [7].
A.4 Qualitative Analysis
A.4.1 Grad-CAM [33] Visualization
Fig. 7 shows the results of class activation maps for SCHEMEformer-PPAA-44-e8-S12 model for a few examples from the validation set of ImageNet-1K dataset. The stronger heatmap responses around the salient features of an object (e.g., body of a bird, cat) shows that the model ignores the background and attends to more discriminative spatial regions. Fig. 7 also shows the qualitative comparison with a few existing methods such as ResNet-50, DeiT-S, Poolformer etc. of similar complexity. It demonstrates that SCHEMEformer-PPAA-44-e8-S12 attends to the complete object class and less spurious features showing that it is better than the competing methods.
A.4.2 Attention Visualization
Figure 8 shows the results of attention maps with and without CCA on Imagenet dataset. We find that there are differences in attention between using or not using CCA. CCA increases the tendency of the model to attend to the regions where the class is present, which is denoted by green ellipses.
(a) Accuracy-FLOPs
(b) Accuracy-Parameters
(c) Accuracy-Throughput
Figure 9: [Zoomed version] Comparison of the proposed SCHEMEformer family, derived from the Metaformer-PPAA-S12 model [55] with higher expansion ratios in the MLP blocks, and many SOTA transformers from the literature. The SCHEMEFormer family establishes a new Pareto frontier (optimal trade-off) for a) accuracy vs. FLOPs, b) accuracy vs model size, and c) accuracy vs, throughput. SCHEMEformer models are particularly effective for the design of fast transformers (throughput between 75 and 150 images/s) with small model size.
Figure 10: [Zoomed version] SCHEME tradeoffs. Top: Accuracy vs FLOPs of various SCHEME models with different MLP configurations. Bottom: SCHEME mixer improves accuracy for fixed throughput or vice-versa for various popular ViT architectures.
Figure 11: [Zoomed version] Impact of CCA (SCHEMEformer-44-e8-S12). Top: Evolution of weight $1-\alpha$ across model layers. Bottom: Class separability of output features $\bf y$ (over 50 random classes of ImageNet-1K validation set) for model trained with and without CCA.